xx ReductionTest

ReductionTest

Christian D. Ott <cott@aei.mpg.de>

April 18 2006

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11domain10Scope: shared from GRIDx01KEYWORD

11

7 Interfaces

General Implements: reductiontest

Inherits: grhydro grid

Grid Variables

7.0.1 PRIVATE GROUPS

11 x Group Names x 10x Variable Names xx01 Details x11x
11centerxofxmassxhelpers10  x01  compact110
11                    10  dMx  x01 dimensions113
11                    10  dMy  x01 distribution11DEFAULT
11                    10  dMz  x01 group type11GF
11                    10  x01  tags11prolongation=”none”
11                    10  x01 timelevels113
11                    10  x01variable type11REAL
11 centerxofxmass 10  x01  compact110
11                    10  Mx  x01 dimensions110
11                    10  My  x01 distribution11CONSTANT
11                    10  Mz  x01 group type11SCALAR
11                    10  Mr  x01 timelevels111
11                    10  x01variable type11REAL

8 Schedule This section lists all the variables which are assigned storage by thorn Carpet/ReductionTest. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:  11x
11centerxofxmass centerxofxmassxhelpers[3]11x

11x  11x

11

Scheduled Functions

CCTKxINITIAL x comxlocal x setup integrands for com calculation x

11x11After: 11whiskyxinitial
11 x 11Language:11c
11 x 11Sync:  11centerxofxmassxhelpers
11 x 11Type: 11function

11

CCTKxEVOL x comxlocal x setup integrands for com calculation x

11x11After: 11molxevolution
11 x 11Language:11c
11 x 11Sync:  11centerxofxmassxhelpers
11 x 11Type: 11function

11

CCTKxPOSTREGRIDINITIAL x comxlocal x setup integrands for com calculation x

11x11Language:11c
11 x 11Sync:  11centerxofxmassxhelpers
11 x 11Type: 11function

CCTKxPOSTREGRID x comxlocal x setup integrands for com calculation x

11x11Language:11c
11 x 11Sync:  11centerxofxmassxhelpers
11 x 11Type: 11function

11

CCTKxANALYSIS x com x com reductions x

11x11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function